OpenStudy (anonymous):

Traveling at 45 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 240 miles from Orlando to Miami? 5 195 285 320

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

Im guessing 5

OpenStudy (flufflepuffle):

Well, divide 240 by 45. It equals 5.333. Rounded to the nearest whole number, it would be 5.

OpenStudy (anonymous):

Well that would mean it would take 5 minutes to travel to Miami?

OpenStudy (texaschic101):

d = rt (distance = rate * time) but since we are looking for time, it is d/r = t d/r = t d = 240 r = 45 240/45 = t 5.33333 = t rounded = 5

OpenStudy (flufflepuffle):

Well, 5 hours. Now, we do 5.333x60.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I have another one, @texaschic101,

OpenStudy (texaschic101):

true ^^ it would be 320 minutes

OpenStudy (anonymous):

Find the value of the expression 4a^4 − 2b^2 + 40 when a = 2 and b = 7. 6 39 74 202

OpenStudy (texaschic101):

matter of subbing 4(2^4) - 2(7^2) + 40 4(16) - 2(49) + 40 64 - 98 + 40 104 - 98 6

OpenStudy (anonymous):

Thank you!! I have another one, but ill try it out myself, ill post another question if i need help, thanks again! :)

OpenStudy (texaschic101):

sure thing :)